# DMAT: An End-to-End Framework for Joint Atmospheric Turbulence Mitigation and Object Detection

### v1

@@ 4.3 Experiment settings
We compare our proposed method, DMAT, with three state-of-the-art AT mitigation methods, TMT [44], DATUM [43], and MAMBA [12], and three object detectors, YOLO11 [36], DETR [6], and LWDETR [8] with various model sizes.
The DATUM and TMT AT-mitigation models were used with pre trained weights. The Mamba model was pre-trained using the ($256\times 256$) training COCO dataset described above. Optimization was performed using the Adam optimizer with an initial learning rate set to 0.0001. The model was trained for a total of 100 epochs without the use of early stopping criteria. These conditions were also used for training the combined system DMAT. A sliding window approach was employed for video processing, wherein a set of temporally adjacent frames centred on the current frame was used as input to predict the target frame. This strategy was uniformly applied during both the training and inference stages. In accordance with the findings reported in [3], a window of ten neighboring frames (four forward frames and five backward frames) was adopted, as it achieved good results but also was able to effectively represent temporal correlations.
Object detectors are fine-tuned by taking the pretrained weights of a network and updating them through training on new data, where the inherent AT distortions have been mitigated by the approaches mentioned above. Upper model layers are often frozen to preserve the representations of low-level visual content. In our training system, however, we update all layers within the object detector (as discussed below).
The implementation was carried out in Python, utilizing the PyTorch framework with CUDA acceleration to ensure computational efficiency. All experiments were conducted on a high-performance computing environment featuring NVIDIA 4090 and A100 GPUs.
@@ 5 Results and discussion
5.1 Performance of object detection
The results presented in Table 2 demonstrate that our proposed DMAT framework outperforms existing methods in object detection under AT conditions. Among two-stage approaches employing the same object detector, MAMBA achieves higher performance than TMT and DATUM. However, compared to our proposed DMAT, its performance can be up to 25% lower when using models of similar size. This highlights the effectiveness of our joint framework, where the AT mitigator generates restored frames with semantically enriched features that are more conducive to accurate object detection.
Figures 5(a), 5(b) and 5(c) present the object detection performance of our DMAT in comparison with a variety of independently and sequentially applied baseline models. The graphs include detection results for the non mitigated ground truth (noAT) and turbulent (AT) inputs, serving as benchmarks for upper-bound and lower-bound performances, respectively. The results from the individual models reveal several trends: (i) detection performance generally improves with increasing model size; (ii) the Mamba-based AT mitigation model yields superior performance relative to DATUM and TMT in all cases.
The right hand side of figures 5(a), 5(b) and 5(c) display the same set of methods evaluated specifically on small objects (defined as those smaller than $32\times 32$ pixels), using mAP[0.50-0.95] as the performance metric. These graphs further illustrate that: (i) for all methods, mAP[0.50-0.95] degrades significantly when limited to small objects, which is expected due to the increased blurring effect of turbulence on smaller targets, even when AT-mitigation is applied; and (ii) the DMAT models maintain a performance advantage over individual methods, consistent with the trends observed across all object sizes.
Fig. 6 shows some example detections. All the shown methods apart from DMAT use the medium sized LW-DETR method (28.2 Million parameters) whereas the DMAT method uses the Medium sized architecture. This shows that object detection is obviously difficult for the distorted input (AT), improves with mitigation, and confirms that DMAT achieves superior performance.
@@ 5.1 Performance of object detection
Table 2: Object detection performance in terms of mAP[0.50-0.95] for AT mitigation methods and detectors across COCO subsets and object sizes. Results in gray indicate ‘no AT’ (i.e., undistorted inputs from COCO images), served as upper-bound performance. The top three results in each column are highlighted in green, orange, and yellow, respectively. ‘AT’ refers to non-mitigated, distorted inputs affected by atmospheric turbulence (simulated using P2S). Parameters column shows number of parameters in the object detector and (where applicable) the additional number of parameters in the AT-mitigator
AT Mitigation Detector Params (M) AllCls-AllSz AllCls-Small Top10-AllSz Top10-Small CarPerson-AllSz CarPerson-Small
no AT YOLO11n 2.6 0.419 0.089 0.279 0.080 0.437 0.313
YOLO11s 9.4 0.509 0.099 0.364 0.095 0.505 0.370
YOLO11m 20.1 0.547 0.122 0.410 0.107 0.547 0.417
YOLO11l 25.3 0.548 0.129 0.430 0.110 0.569 0.430
YOLO11x 56.9 0.576 0.133 0.450 0.114 0.590 0.451
DETR-Med 41.3 0.483 0.307 0.312 0.204 0.448 0.309
DETR-Large 60.0 0.525 0.363 0.346 0.234 0.468 0.468
LWDETR-Tiny 12.1 0.523 0.410 0.384 0.273 0.505 0.375
LWDETR-Small 14.6 0.561 0.431 0.428 0.313 0.534 0.400
LWDETR-Med 28.2 0.628 0.480 0.464 0.343 0.566 0.427
AT YOLO11n 2.6 0.050 0.010 0.020 0.004 0.062 0.014
YOLO11s 9.4 0.090 0.014 0.040 0.004 0.099 0.020
YOLO11m 20.1 0.093 0.012 0.040 0.005 0.104 0.022
YOLO11l 25.3 0.101 0.010 0.040 0.005 0.114 0.022
YOLO11x 56.9 0.124 0.029 0.060 0.009 0.148 0.034
DETR-Med 41.3 0.092 0.035 0.031 0.003 0.092 0.021
DETR-Large 60.0 0.066 0.026 0.035 0.008 0.104 0.025
LWDETR-Tiny 12.1 0.158 0.068 0.057 0.000 0.136 0.039
LWDETR-Small 14.6 0.149 0.073 0.070 0.014 0.147 0.049
LWDETR-Med 28.2 0.192 0.061 0.081 0.017 0.164 0.049
DATUM YOLO11n 5.8+2.6 0.128 0.034 0.110 0.026 0.217 0.097
YOLO11s 5.8+9.4 0.300 0.054 0.134 0.038 0.259 0.117
YOLO11m 5.8+20.1 0.307 0.057 0.151 0.033 0.276 0.130
YOLO11l 5.8+25.3 0.301 0.065 0.151 0.035 0.289 0.131
YOLO11x 5.8+56.9 0.317 0.060 0.166 0.032 0.291 0.132
DETR-Med 5.8+41.3 0.218 0.076 0.108 0.034 0.195 0.074
DETR-Large 5.8+60.0 0.243 0.100 0.120 0.051 0.216 0.087
LWDETR-Tiny 5.8+12.1 0.293 0.158 0.190 0.074 0.265 0.120
LWDETR-Small 5.8+14.6 0.332 0.167 0.209 0.096 0.289 0.134
LWDETR-Med 5.8+28.2 0.369 0.200 0.231 0.098 0.306 0.140
TMT YOLO11n 22.9+2.6 0.229 0.029 0.107 0.021 0.203 0.082
YOLO11s 22.9+9.4 0.294 0.028 0.149 0.024 0.241 0.094
YOLO11m 22.9+20.1 0.276 0.042 0.154 0.028 0.242 0.092
YOLO11l 22.9+25.3 0.301 0.054 0.161 0.028 0.263 0.111
YOLO11x 22.9+56.9 0.307 0.038 0.168 0.026 0.256 0.092
DETR-Med 22.9+41.3 0.204 0.062 0.101 0.029 0.162 0.045
DETR-Large 22.9+60.0 0.224 0.087 0.129 0.040 0.178 0.052
LWDETR-Tiny 22.9+12.1 0.297 0.134 0.188 0.074 0.237 0.097
LWDETR-Small 22.9+14.6 0.301 0.143 0.196 0.076 0.257 0.106
LWDETR-Med 22.9+28.2 0.344 0.165 0.226 0.088 0.271 0.119
MAMBA YOLO11n 2.8+2.6 0.249 0.024 0.120 0.026 0.228 0.098
YOLO11s 2.8+9.4 0.310 0.038 0.176 0.028 0.267 0.117
YOLO11m 2.8+20.1 0.312 0.046 0.177 0.030 0.285 0.130
YOLO11l 2.8+25.3 0.337 0.055 0.179 0.031 0.294 0.135
YOLO11x 2.8+56.9 0.350 0.037 0.191 0.031 0.292 0.128
DETR-Med 2.8+41.3 0.227 0.073 0.113 0.031 0.200 0.067
DETR-Large 2.8+60.0 0.275 0.097 0.135 0.041 0.217 0.076
LWDETR-Tiny 2.8+12.1 0.316 0.164 0.193 0.076 0.270 0.115
LWDETR-Small 2.8+14.6 0.354 0.161 0.224 0.094 0.295 0.136
LWDETR-Med 2.8+28.2 0.364 0.186 0.249 0.105 0.310 0.138
DMAT-Tiny (ours) 2.8+12.1 0.374 0.188 0.214 0.108 0.348 0.178
DMAT-Small (ours) 2.8+14.6 0.407 0.234 0.276 0.134 0.373 0.194
DMAT-Med (ours) 2.8+28.2 0.419 0.234 0.282 0.137 0.385 0.199
Table 3: Comparison of AT Mitigation methods using Mean PSNR, SSIM, and LPIPS metrics across three datasets: All, CarPerson, and Top10 categories. Best values are highlighted in green, second in orange, third in yellow.
AT Mitigation All CarPerson Top10
@@
AT (no mitigation) 20.946 0.512 0.546 20.210 0.488 0.564 20.668 0.513 0.550
DATUM 23.092 0.644 0.356 22.390 0.632 0.360 22.910 0.649 0.343
TMT 23.332 0.654 0.348 22.592 0.641 0.350 23.167 0.661 0.331
MAMBA 23.763 0.667 0.378 23.084 0.657 0.380 23.693 0.677 0.360
DMAT (ours) 23.841 0.671 0.373 23.220 0.663 0.376 23.861 0.683 0.357
Figure 6: Comparison of AT-mitigation and Object Detection Methods. All methods apart from DMAT use the medium sized LW-DETR method (28.2 Million parameters) whereas the DMAT method uses the Medium sized architecture.
5.2 Performance of AT mitigation
@@ 6 Conclusion
In this paper, we introduce the first end-to-end framework that jointly addresses AT mitigation and object detection. By tightly coupling a 3D Mamba-based turbulence mitigation module with a transformer-based object detector, the proposed architecture enables mutual enhancement between low-level restoration features and high-level semantic representations. This integration not only preserves crucial object-level details often lost during restoration but also significantly improves detection accuracy under severe atmospheric distortions. Extensive experiments on synthetically generated turbulent datasets derived from the COCO benchmark demonstrated that our combined system outperforms both individually trained AT-mitigation and object detection pipelines. Notably, the unified architecture yielded up to a 15% improvement in mAP[0.50-0.95] across various detector types and turbulence conditions, and maintained strong performance even on challenging subsets such as small object categories and surveillance-focused object classes. These results highlight the advantages of bidirectional feature flow in restoring spatial-temporal consistency and enhancing detection robustness. The framework further shows that alternating loss function optimization—rather than weighted joint loss—achieves better convergence and output quality, mitigating the issue of visual degradation commonly observed when attempting to balance restoration and detection objectives. Although the combined architecture proposed gives superior performance compared to the non combined models and methods it does have a larger memory requirement due to the models needing to be loaded into memory together during training.

### v3

@@ 4.2 Experiment settings
We compare our proposed DMAT with three state-of-the-art AT mitigation methods: TMT [40], DATUM [39], and MAMBA [12], as well as three object detectors: YOLOv11 [19], DETR [5], and LW-DETR [8], each evaluated at multiple model sizes. Note that we do not apply joint optimization to the baseline methods, as doing so would incorporate our contributions into their architectures, resulting in an unfair comparison.
All models are fine-tuned on the new synthetic datasets described above. All models use the Adam optimizer with a learning rate of 0.0001 for 100 epochs, without early stopping. A sliding window approach was used during both training and inference, where each target frame was predicted from ten temporally adjacent frames (five preceding, four following), following the setup in [3], which balances performance and temporal context. The implementation was carried out in Python, utilizing the PyTorch framework with CUDA acceleration to ensure computational efficiency. All experiments were conducted on NVIDIA 4090 and A100 GPUs.
@@ 5 Results and discussion
We evaluate system performance in both object detection and atmospheric turbulence (AT) mitigation using our proposed synthetic datasets and real AT videos. For synthetic datasets, where ground truth is available, we report quantitative metrics, while real AT videos allow only for subjective visual assessment. Object detection performance is measured using Average Precision (AP) across multiple Intersection-over-Union (IoU) thresholds (0.50 to 0.95 in 0.05 increments, denoted AP[0.50–0.95]). We report Mean Average Precision (mAP), which averages AP across all object categories. For AT mitigation, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [35], and the perceptual similarity metric LPIPS [38].
(a) Static scenes with car and Person
(b) Static scenes with top 10 objects
(c) Dynamic scenes with car and Person
@@
5.1 Synthetic datasets
5.1.1 Performance of object detection.
The results presented in Table 1 demonstrate that our proposed DMAT framework outperforms existing methods in object detection under AT conditions on COCO datasets. Among two-stage approaches employing the same object detector, MAMBA achieves higher performance than TMT and DATUM. However, compared to our proposed DMAT, its performance can be up to 25% lower when using models of similar size. This highlights the effectiveness of our joint framework, where the AT mitigator generates restored frames with semantically enriched features that are more conducive to accurate object detection.
Fig. 5 shows the object detection performance of our DMAT model compared with various baseline models, applied either independently or sequentially, on the COCO and GOT-10k datasets. The graphs include results for non-mitigated ground truth inputs (noAT) and turbulent inputs (AT), serving as upper- and lower-bound benchmarks, respectively. We present results for 2 and 10 object classes; performance across all object sizes and classes follows a similar trend, as shown in Fig. 1. The results highlight two key trends: (i) detection performance generally improves with increasing model size; and (ii) the Mamba-based AT mitigation model consistently outperforms both DATUM and TMT across all cases.
The right hand side of Fig. 5 display the same set of methods evaluated specifically on small objects (defined as those smaller than $32\times 32$ pixels). These graphs further illustrate that: (i) for all methods, mAP[0.50-0.95] degrades significantly when limited to small objects, which is expected due to the increased blurring effect of turbulence on smaller targets, even when AT-mitigation is applied; and (ii) the DMAT models maintain a performance advantage over individual methods, consistent with the trends observed across all object sizes. Fig. 6 shows some example detections. All the shown methods apart from DMAT use the medium sized LW-DETR method (28.2 Million parameters) whereas the DMAT method uses the Medium sized architecture. This shows that object detection is obviously difficult for the distorted input (AT), improves with mitigation, and confirms that DMAT achieves superior performance.
AT Mitigation Static scenes Dynamic scenes
All Top10 CarPerson CarPerson
@@ 5.1 Synthetic datasets
AT 20.946 0.512 0.546 20.668 0.513 0.550 20.210 0.488 0.564 22.252 0.591 0.466
DATUM 23.092 0.644 0.356 22.910 0.649 0.343 22.390 0.632 0.360 22.564 0.693 0.373
TMT 23.332 0.654 0.348 23.167 0.661 0.331 22.592 0.641 0.350 22.747 0.712 0.367
MAMBA 23.763 0.667 0.378 23.693 0.677 0.360 23.084 0.657 0.380 23.297 0.783 0.379
DMAT (ours) 23.841 0.671 0.373 23.861 0.683 0.357 23.220 0.663 0.376 23.312 0.831 0.371
Table 2: Comparison of AT Mitigation methods for synthetic static scenes across two datasets: All and Top10 categories. Best values are highlighted in bold, underline, and italic.
Figure 6: Comparison of AT-mitigation and Object Detection Methods. All methods apart from DMAT use the medium sized LW-DETR method whereas the DMAT method uses the Medium sized architecture.
5.1.2 Performance of AT mitigation.
@@
5.2 Real AT sequences
Figure 8: Qualitative detection results on real-world atmospheric turbulence videos from the CLEAR dataset. From left to right: distorted input, BasicVSR++, DATUM, and our proposed DMAT. All baseline methods adopt the medium-sized LW-DETR method, while DMAT uses its own medium-sized configuration. Zoom in for better visualization.
To further verify the generalisation ability of our proposed DMAT framework in real atmospheric turbulence scenarios, we tested it on real turbulent videos from the CLEAR dataset [2]. These videos have no ground truth labels, so we can only make visual comparisons. Fig. 8 shows a qualitative comparison between DMAT and two typical restoration methods: BasicVSR++ [7] and DATUM [39]. To keep the comparison fair, all baseline methods used the medium-size LW-DETR detector, while DMAT used its own medium-size version.
In the van scene (top row), where the van moves from left to right, the raw input has strong geometric distortion and edge blur, leading to very low detection confidence. BasicVSR++ reduces some blur but still gives distorted object shapes and unstable detection. DATUM gives clearer object outlines than BasicVSR++, but there are still turbulence artefacts around the edges, which affect detection. In contrast, DMAT removes most distortions, restores sharper and more complete vehicle structures (like wheels and windows), and gives much higher detection confidence.
In the human scene (bottom row), where three men are moving objects in their hands, the raw input shows flickering and motion ghosting, so the detector often misses the targets. BasicVSR++ partly restores the human shapes, but detection is still unstable. DATUM detects several targets but with low confidence scores. In contrast, DMAT restores clear and complete human shapes, allowing the detector to find all targets with high confidence. This shows that DMAT generalises well from synthetic data to real turbulent conditions.
Overall, DMAT can keep object structures clear and consistent under real turbulence, while traditional restoration-only methods cannot achieve this.
5.3 Ablation study
Experiment 1: As seen in Table 1 and Fig. 5, when the AT mitigator is removed, the performance of object detection drops by approximately 50% on average, with an even larger drop for smaller objects.
Experiment 2: When the object detection module is removed, the performance of AT mitigation drops by up to approximately 1.5% in term of PSNR.
Experiment 3: When the feature sharing between the two modules is removed (red lines in Fig. 3), the performance degradation is reduced by up to 1% in terms of PSNR and 2% in terms of mAP, indicating that better integration of features contributes to both improved reconstruction quality and detection accuracy.
@@ 6 Conclusion
We propose the first end-to-end framework that jointly tackles atmospheric turbulence (AT) mitigation and object detection. By integrating a 3D Mamba-based restoration module with a transformer-based detector, our architecture enables mutual enhancement between low-level and high-level features. This improves detection accuracy under severe distortions while preserving object-level details. Experiments on turbulent COCO datasets show our method outperforms separate AT-mitigation and detection models, achieving up to a 15% mAP[0.50–0.95] gain. Bidirectional feature flow improves robustness, especially for small and surveillance-relevant objects. While performance is strong, the joint model requires more memory during training due to its combined architecture.